OpenStudy (goformit100):

I know a way by which i can make i can get total of 120 by using five zeros 0,0,0,0,0 and any one mathematical operator. Do you ?

OpenStudy (goformit100):

@TessaMarie08181997 @Koikkara

OpenStudy (goformit100):

@.Sam.

OpenStudy (anonymous):

enlighten me.

OpenStudy (anonymous):

you have to use the operator called factorial

OpenStudy (koikkara):

@goformit100 ...ask tough questions..these r just simple...lol

OpenStudy (koikkara):

@brookie5150 ...well,Here we have used operator called " factorial ". As you know that 0! = 1 so, = ( 0! + 0! + 0! + 0! + 0! ) ! = ( 1 + 1 + 1 + 1 + 1 ) ! = (5 )! = ?......u got to answer....n be the heroin...

sam (.sam.):

120...

OpenStudy (koikkara):

@goformit100 @brookie5150 ...she knows....lol

sam (.sam.):

5 zeros 200-60-10-10=...